The perimeter of a rectangle is 52. The length is 6 times greater than the width. Find the length and width of this rectangle.
Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

The perimeter of a rectangle is 52. The length is 6 times greater than the width. Find the length and width of this rectangle.
------------
2L + 2W = 52
L + W = 26
L = W + 6W = 7W
L + W = 26
7W + W = 26
8W = 26
W = 13/4
L = 91/4

Answer by solver91311(24713)



Let represent the width. Then the length must be .

The perimeter of a rectangle is given by:



Substitute what you know:



And solve for , then calculate for the length.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism